OpenStudy (anonymous):

find |vector d + vector e| given that |vector d|=3 and |vector e| = 5 where the angle between these two vector is 50 degrees.

OpenStudy (anonymous):

is it 18.4

OpenStudy (anonymous):

the answer is 7.30 :( but i dont knw how to get it

OpenStudy (anonymous):

nw mine is cmng 13.9

OpenStudy (anonymous):

what did you do ?

OpenStudy (anonymous):

i did |d+e|^2=|d|^2+|e|^2 +2|d||e|cos |d+e|^2=25+9+30*50pi/180 |d+e|^2=34+5.23 =39.23 |d+e|^2=39.23 |d+e|=6.2

OpenStudy (anonymous):

i got it :) ! you just caclulated the cos value wronge :) ! but thanks alot i got the idea from your solution :) !